OpenStudy (anonymous):

A pyramid has a square base with sides 8" and a slant height of 5". Total Area =

OpenStudy (agent47):

Area of the base=8*8=64

OpenStudy (agent47):

Area of the side: (1/2)b*h=(1/2)*8*h=4h

OpenStudy (agent47):

h=5 A of the side = 4*5=20

OpenStudy (agent47):

Total area: 64+4*20=144